Institute for Astronomy Home
IFA Home Page   |    Search   |    Other Editions    No. 42 - 2012 
  All Articles  


UH Telescope Upgrades Delayed by Drive Problem

by Colin Aspin, Director, UH 2.2-meter telescope

telescope mirror

The mirror of the UH 2.2-meter telescope shows the reflections of Greg Osterman (right) and Colin Aspin.

The University of Hawai‘i 2.2-meter telescope has been located on the volcanic peak of Mauna Kea on the Big Island of Hawai‘i for over 40 years. It was the first large telescope on the mountain and has produced many exciting discoveries, including the detection of the first outer solar system Kuiper Belt object. During the last two decades, however, it has been overshadowed by the newer, larger telescopes on Mauna Kea. But even over this period, the UH88, as it is colloquially called (it was originally named the UH 88-inch telescope), has remained the workhorse for astronomical research for the faculty and graduate students at the Institute for Astronomy. For while IfA astronomers have access to all the larger telescopes on Mauna Kea, the 2.2-meter telescope is the only professional telescope on Mauna Kea that is totally controlled by UH. Now, however, its age has become increasingly detrimental to the efficiency and productivity of the astronomers using it.

In the mid-2000s, IfA developed a plan to replace the UH88 with either a modern survey telescope (such as Pan-STARRS) or a larger telescope. However, for many reasons, including the economic downturn of the late 2000s, these plans have not yet become a reality, so we are continuing our efforts keep the UH88 operational and scientifically productive for at least another five years. This is proving to be a rather challenging task requiring significant effort and ingenuity from the small, dedicated team of engineers and scientists working on it. We have started around 20 upgrade projects aimed at improving the reliability and efficiency of the telescope and instrumentation.

Unfortunately, these upgrades are now on hold while we solve a more serious problem—a fault with the telescope hour angle drive, which enables the telescope to track astronomical objects as they rise and set. Recently, this drive has begun to fail to complete movements and to stop while following astronomical objects. The components of the drive are all over 40 years old, and our engineering team has been studying their design and operation while attempting to track down the source of the failure.

The detective work involved in tracking down the cause of this problem has so far involved checking five different systems. First, we looked at the high-level computer programming to understand what event triggers the error that stops the telescope motion. Next, we examined the motion control electronics, that is, the signals between the telescope control computer and the telescope itself. The third system checked was the mechanics. We confirmed that the gears, shafts, bearings, and brakes forming the hour angle drive operate correctly. Fourthly, we checked the hydraulic pump and pressure lines to make sure they are working and leak-free. Having ruled out problems with those systems, we finally examined the oil viscosity.

The telescope has a hydrostatic oil bearing that is supposed to have a thickness of 8 thousandths of one inch for reliable operation. The oil used has to possess very specific viscosity characteristics over the operating temperature of the oil, which is 32° to 68° F (0° to 20° C). The oil originally used in the hydrostatic bearing was called Mobil “Flying Horse" oil, but unfortunately the company terminated production around 15 years ago. A replacement oil was installed around five years ago, and this was deemed the best possible substitute then available. However, after much investigation, it is seeming more and more likely that either the viscosity characteristics of that oil are not close enough to those of the original oil, or that the oil has become contaminated over time. After an in-house study of oil characteristics, 150 gallons (nearly 600 liters) of a more modern synthetic oil have been ordered, and we hope that this will solve our current dilemma.

Once the problem with the hour angle drive is resolved, and normal observing has been resumed, we will again turn our attention to our upgrade projects. More on those in the next issue of this newsletter.